Pre Mustandes process. co m with two de Mixing Department Bling Dawn During May. Priemerated $130.000 in producing 40.000 (10 could you add work as well please
Answer #1

Question 1

Part 1

Unit cost per gallon of mustard processed in the mixing department in May = (direct materials + direct labor + manufacturing overhead)/ Total gallons = (16000+5600+11400)/40000 = $0.825

Part 2

unit cost per case of mustard incurred by the Bottling Department in May =(direct materials + direct labor + manufacturing overhead)/ Total cases =(6,000 + 3,900 + 8,100)/10,000 = $1.80

Part 3

Total cost of the 10,000 cases transferred to finished goods in May = (0.825 * 4 gallons/case) + 1.80 = 5.10 * 10,000 = $51,000

Part 4

Amount charged to cost of goods sold in May = 9,500 * 5.10 = $48450

Question 2

Part 5

Contribution margin ratio = contribution margin per unit / selling price per unit

Contribution margin per unit = selling price per unit – variable cost per unit = 50 – (70%*50) = 50-35 = $15

Contribution margin ratio = 15/50= 30%

Part 6

Breakeven in units = Fixed cost / Contribution margin per unit   = 7500/15 = 500 units

Part 7

Desired sales in dollars = (fixed cost + targeted operating income)/ Contribution margin per unit (7500+5000)/15 = 833 units

Desired sales in dollars = 833*50 = $41650

Part 8

Contribution margin ratio = contribution margin per unit / selling price per unit

Contribution margin per unit = selling price per unit – variable cost per unit = 55 – (70%*55) = 55-38.5 = $16.5

Contribution margin ratio = 16.5/55= 30%